# On degenerate metric resolvent: well-definedness and new convergence analysis

### v3

@@ 5.2 Discussions and related works
First, Lemma 5.1 makes the connection between the degenerate PPA (2) and the original inclusion problem (1). One can always show the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$ without Lemma 5.1. This is Lemma 5.1 that tells us what the degenerate PPA (2) solves for is actually a point of $\mathsf{zer}\mathcal{A}$ projected onto the range space of degenerate metric.
Assumption 4—a stronger condition than Assumption 3—is not needed in Theorem 5.3, and thus, $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is also not necessarily maximally monotone here. The single-valued conditions listed in Theorem 4.2 are also not needed. In other words, $\mathcal{T}$ is allowed to be multi-valued, since we are only concerned with $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, which is already well-defined by Assumptions 1 and 3. This is fundamentally different from [22, Theorem 3.4], where $\mathcal{T}$ has to be well-defined.
[22, Theorem 3.4] discussed the convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in the finite-dimensional setting, where the proof heavily relied on the assumption of continuity of $\mathcal{T}$. We here argue that the continuity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ would suffice for [22, Theorem 3.4], which is obviously true in finite-dimensional case by Fact 2. However, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$ is not weakly continuous generally in infinite-dimensional case. Instead, we exploited the weak-to-strong continuity of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ in Lemma 2.8.
@@
Our proof uses the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$. An alternative way is to use the maximal monotonicity of $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ (i.e., Assumption 3) to finish the proof as well, see Appendix B.
Finally, we stress that $\{x^{k}\}_{k\in\mathbb{N}}$ may be unbounded due to the uncontrolled component of $x_{\mathsf{k}}^{k}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x^{k}$. Thus, $\{x^{k}\}_{k\in\mathbb{N}}$ may not even possess any weak sequential cluster point, and there is no conclusion about the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$. The only exception occurs in the non-degenerate case, where the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ is obviously equivalent to that of $\{x^{k}\}_{k\in\mathbb{N}}$. Assumption 3 in this case boils down to the usual maximal monotonicity of $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$, as stated in Corollary 3.7.
@@ 5.3 Convergence in the whole space
To proceed further, if $\mathcal{T}$ is well-defined, it is then natural to ask when the iterations (2) converge in the whole space. At least, the boundedness of $\{x^{k}\}_{k\in\mathbb{N}}$ should be a basic requirement. To this end, [10, Theorem 2.9] assumed the Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$. Here, we propose the equivalent conditions as
@@
For any two inputs $x_{i,\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$ for $i=1$ and 2, and the corresponding $y_{i,\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{i,\mathsf{r}})$,
@@
Under Assumptions 1, 3 and the same setting of Theorem 4.2, then the following 3 conditions are equivalent:
@@
(iii) Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$ [10, Theorem 2.9].
The proof is postponed in Appendix C. Indeed, all the conditions in Proposition 5.5 are equivalent to $\|\mathcal{T}x_{1}-\mathcal{T}x_{2}\|\leq\xi\|x_{1}-x_{2}\|_{\mathcal{Q}}$, which can be found in [32, Theorem 3.3] and [11, Lemma 3.2].
@@
Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2), which satisfies Assumptions 1, 3, 5 and any condition of Theorem 4.2 and Proposition 5.5. Furthermore, if any of the following conditions hold:
(i) every weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ lies in $\mathsf{Fix}\mathcal{T}$;
@@
then $\{x^{k}\}_{k\in\mathbb{N}}$ weakly converges to some $x^{\star}\in\mathsf{zer}\mathcal{A}$, as $k\rightarrow\infty$.
@@
Proposition 5.5 guarantees the boundedness of $\{x^{k}\}_{k\in\mathbb{N}}$. Then there exist two subsequences $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ and $\{x^{k_{j}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v_{1}^{*}$ and $x^{k_{j}}\rightharpoonup v_{2}^{*}$. By Step-3 of the proof of Theorem 5.3, we have $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}v_{1}^{*}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}v_{2}^{*}$ and further $\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}$, i.e., all the weak sequential cluster points of $\{x^{k}\}_{k\in\mathbb{N}}$ share the same projection onto $\mathsf{ran}\mathcal{Q}$.
Condition (i): it implies that $v_{1}^{*}=\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}=v_{2}^{*}$. This shows that the weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ is unique, and lies in $\mathsf{Fix}\mathcal{T}=\mathsf{zer}\mathcal{A}$. The weak convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ is obtained by [4, Lemma 2.46].
Condition (ii): Assumption 4-(i) implies that $\mathsf{gra}\mathcal{A}$ is closed in $\mathcal{H}^{\text{weak}}\times\mathcal{H}^{\text{strong}}$. Take a subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v^{*}$. Then $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$, since $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$. Considering $\mathcal{A}\mathcal{T}x^{k_{i}}\owns\mathcal{Q}(x^{k_{i}}-\mathcal{T}x^{k_{i}})$, combining with $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$ and $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$, it then follows from the closedness of $\mathsf{gra}\mathcal{A}$ that $v^{\star}\in\mathsf{zer}\mathcal{A}=\mathsf{Fix}\mathcal{T}$, which leads to Condition (i).
Condition (iii): Considering the subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ in (ii), the demiclosedness of $\mathcal{I}-\mathcal{T}$ yields that $(\mathcal{I}-\mathcal{T})v^{*}=0$, i.e., $v^{*}\in\mathsf{Fix}\mathcal{T}$, which also leads to Condition (i).
Conditions (i) and (ii) correspond to [10, Theorem 2.9, Corollary 2.10], respectively. Condition (iii), as mentioned in [10, Remark 2.11], is hard to control. The key problem here is that $\mathcal{T}$ is not necessarily nonexpansive, and thus the general demiclosedness of $\mathcal{I}-\mathcal{T}$ cannot be obtained by Browder’s principle [4, Theorem 4.27]. However, it can be tackled case-by-case, by exploiting particular structures of $\mathcal{A}$ and $\mathcal{Q}$, and additional assumptions. For instance, [11, Lemma 3.4], in a context of a specific class of preconditioned ADMM, proposed several additional assumptions (e.g., compactness, weak closedness) to make $\mathcal{I}-\mathcal{T}$ demiclosed.
[16, Theorem 3.3] seems the first attempt to establish the whole convergence under degenerate metric in a finite-dimensional setting, based on the continuity of proximity operator. This technique, together with [22, Theorem 3.4], is not applicable for infinite-dimensional setting, since $\mathcal{T}$ is generally not weakly continuous.
@@ 6.1 Douglas–Rachford splitting (DRS)
The scheme of DRS has been given as (3), with the degenerate PPA interpretation of (4). The well-definedness of the DRS is obvious from (3). It can also be verified from (4) only via Assumption 4. First, it is clear that $\mathcal{A}:\mathcal{U}^{3}\mapsto 2^{\mathcal{U}}\times 2^{\mathcal{U}}\times\mathcal{U}$ is maximally monotone. Second, observe that $\mathsf{ran}\mathcal{A}=\mathcal{U}^{3}=\mathcal{H}$ and $0\in\mathsf{ran}\mathcal{Q}$, and thus, $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$, which fulfills Assumption 4-(ii). Then it follows that the scheme (3) admits any $(u,w,z)\in\mathcal{U}^{3}$ in any iterative step $k$.
A standard result in the seminal work [23] claims $z^{k}\rightharpoonup z^{\star}$ as $k\rightarrow\infty$, if $\tau>0$. This can also easily be obtained by our Theorem 5.3. The weak convergence of the solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ had been a longstanding problem until the advent of [30]. To obtain this result, [30] used some sophisticated tools, e.g., quasi-Fejér convergence, strong/weak topologies. In [33, Theorem 4.2], the author used many variants of PPA forms of DRS (especially the complicated prediction+correction framework initialized by [21, 20, 19]). Now, the next Proposition 6.1 shows that the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ can be directly obtained from (4) alone and via our Theorem 5.6. This is an improvement, compared to [33, Remark 3.1-(iii)], which once claimed that the convergences of $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ remain unclear from the PPA form of (4) only. This demonstrates the great potential of using the PPA interpretation. In addition, similar to [30], Proposition 6.1 also removes the assumption of maximal monotonicity of $\partial f+\partial g$, but without using the instrumental result of [3, Corollary 3] as [30] did.
@@
Let $\{u^{k},w^{k},z^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (3). If $\tau>0$, then $(u^{k},w^{k},z^{k})\rightharpoonup(u^{\star},w^{\star},z^{\star})$ as $k\rightarrow\infty$, where $u^{\star}$ is an optimal solution to $\min_{u}f(u)+g(u)$.
@@
Putting $b_{2}=b_{1}-a+c$ (indicated by the last line) into the first line, we obtain $\tau\partial g(b_{1})+b_{1}-a+c-c=\tau\partial g(b_{1})+b_{1}-a\owns 0$, leading to $b_{1}=\operatorname{prox}_{\tau g}(a)$, and thus, $b_{2}=\operatorname{prox}_{\tau g}(a)-a+c=c-\operatorname{prox}_{g^{*}/\tau}(a)$. $b$ is obviously single-valued, which meets Theorem 4.2. Then, from the well-known nonexpansiveness of proximity operator (see [15, Lemma 2.4] for example), we have
@@
$=$ $\big{\|}\operatorname{prox}_{\tau g}(a)-\operatorname{prox}_{\tau g}(a^{\prime})\big{\|}+\big{\|}c-\operatorname{prox}_{g^{*}/\tau}(a)-c^{\prime}+\operatorname{prox}_{g^{*}/\tau}(a^{\prime})\big{\|}$

### v4

@@ 5.2 Discussions and related works
First, Lemma 5.1 makes the connection between the degenerate PPA (2) and the original inclusion problem (1). One can always show the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in $\mathrm{Fix}(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})$ without Lemma 5.1. This is Lemma 5.1 that tells us what the degenerate PPA (2) solves for is actually a point of $\mathrm{zer}\mathcal{A}$ projected onto the range space of degenerate preconditioner.
Assumption 4—a stronger condition than Assumption 3—is not needed in Theorem 5.3, and thus, it is not necessary to assume $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ to be maximally monotone here. Moreover, Theorem 4.2 is not needed. In other words, $\mathcal{T}$ is allowed to be multi-valued, since we are only concerned with $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ here, which is already well-defined by Assumptions 1 and 3. This is fundamentally different from [33, Theorem 3.4], where $\mathcal{T}$ has to be well-defined.
[33, Theorem 3.4] discussed the convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in the finite-dimensional setting, where the proof heavily relied on the assumption of continuity of $\mathcal{T}$. We here argue that the continuity of $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}:\mathrm{ran}\mathcal{Q}\mapsto 2^{\mathrm{ran}\mathcal{Q}}$ would suffice for [33, Theorem 3.4], and this is obviously true in finite-dimensional case by Fact 2. However, $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}$ is not weakly continuous generally in infinite-dimensional case. Instead, we exploited the weak-to-strong continuity of $\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}$ in Lemma 2.9.
@@
Our proof uses the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathrm{ran}\mathcal{Q}}$. An alternative way is to use the maximal monotonicity of $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathrm{ran}\mathcal{Q}}$ (i.e., Assumption 3) to finish the proof as well, see Appendix B.
Finally, we stress that $\{x^{k}\}_{k\in\mathbb{N}}$ may be unbounded due to the uncontrolled component of $x_{\mathsf{k}}^{k}:=\mathcal{P}_{\mathrm{ker}\mathcal{Q}}x^{k}$. Thus, $\{x^{k}\}_{k\in\mathbb{N}}$ may not even possess any weak sequential cluster point, and there is no conclusion about the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$. The only exception occurs in the non-degenerate case, where the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ is obviously equivalent to that of $\{x^{k}\}_{k\in\mathbb{N}}$.
@@ 5.3 Convergence in the whole space
To proceed further, if $\mathcal{T}$ is well-defined, it is then natural to ask when the iterations (2) converge in the whole space. At least, the boundedness of $\{x^{k}\}_{k\in\mathbb{N}}$ should be a basic requirement. To this end, [13, Theorem 2.9] assumed the Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$. Here, we propose the equivalent conditions as
@@
For any two inputs $x_{i,\mathsf{r}}\in\mathrm{ran}\mathcal{Q}$ for $i=1$ and 2, and the corresponding $y_{i,\mathsf{r}}=(\mathcal{P}_{\mathrm{ran}\mathcal{Q}}\circ\mathcal{T})(x_{i,\mathsf{r}})$,
@@
Under Assumptions 1, 3 and the same setting of Theorem 4.2, then the following three conditions are equivalent:
@@
(iii) Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$ [13, Theorem 2.9].
The proof is postponed in Appendix C. Indeed, all the conditions here are equivalent to $\|\mathcal{T}x_{1}-\mathcal{T}x_{2}\|\leq\xi\|x_{1}-x_{2}\|_{\mathcal{Q}}$, which can be found in [50, Theorem 3.3] and [15, Lemma 3.2].
@@
Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2), which satisfies Assumptions 1, 3 and any condition of Theorem 4.2 and Proposition 5.5. Furthermore, if any of the following conditions holds:
(i) every weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}\mathcal{T}$;
@@
then $\{x^{k}\}_{k\in\mathbb{N}}$ weakly converges to some $x^{\star}\in\mathrm{zer}\mathcal{A}$, as $k\rightarrow\infty$.
@@
Proposition 5.5 guarantees the boundedness of $\{x^{k}\}_{k\in\mathbb{N}}$. Then there exist two subsequences $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ and $\{x^{k_{j}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v_{1}^{*}$ and $x^{k_{j}}\rightharpoonup v_{2}^{*}$. By Step-3 of the proof of Theorem 5.3, we have $\mathcal{P}_{\mathrm{ran}\mathcal{Q}}v_{1}^{*}=\mathcal{P}_{\mathrm{ran}\mathcal{Q}}v_{2}^{*}$ and further $\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}$, i.e., all the weak sequential cluster points of $\{x^{k}\}_{k\in\mathbb{N}}$ share the same projection onto $\mathrm{ran}\mathcal{Q}$.
Condition (i): it implies that $v_{1}^{*}=\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}=v_{2}^{*}$. This shows that the weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ is unique, and lies in $\mathrm{Fix}\mathcal{T}=\mathrm{zer}\mathcal{A}$. The weak convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ is obtained by [5, Lemma 2.46].
Condition (ii): Assumption 4-(i) implies that $\mathrm{gra}\mathcal{A}$ is closed in $\mathcal{H}^{\text{weak}}\times\mathcal{H}^{\text{strong}}$. Take a subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v^{*}$. Then $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$, since $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$. Considering $\mathcal{A}\mathcal{T}x^{k_{i}}\owns\mathcal{Q}(x^{k_{i}}-\mathcal{T}x^{k_{i}})$, combining with $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$ and $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$, it then follows from the closedness of $\mathrm{gra}\mathcal{A}$ that $v^{\star}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}\mathcal{T}$, which leads to Condition (i).
Condition (iii): Considering the subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ in (ii), the demiclosedness of $\mathcal{I}-\mathcal{T}$ yields that $(\mathcal{I}-\mathcal{T})v^{*}=0$, i.e., $v^{*}\in\mathrm{Fix}\mathcal{T}$, which also leads to Condition (i).
Conditions (i) and (ii) correspond to [13, Theorem 2.9, Corollary 2.10], respectively. Condition (iii), as mentioned in [13, Remark 2.11], is hard to control and examine. The key problem here is that $\mathcal{T}$ is not necessarily nonexpansive, and thus the general demiclosedness of $\mathcal{I}-\mathcal{T}$ cannot be obtained by Browder’s principle [5, Theorem 4.27]. However, it can be tackled case-by-case, by exploiting particular structures of $\mathcal{A}$ and $\mathcal{Q}$, and additional assumptions, see, for instance, [15, Lemma 3.4].
@@ 6.1 Douglas–Rachford splitting (DRS)
The scheme of DRS has been given as (3), with the degenerate PPA interpretation of (4). The well-definedness of the DRS is obvious from (3). It can also be verified from (4) only via Assumption 4. First, it is clear that $\mathcal{A}:\mathcal{U}^{3}\mapsto 2^{\mathcal{U}}\times 2^{\mathcal{U}}\times\mathcal{U}$ is maximally monotone. Second, observe that $\mathrm{ran}\mathcal{A}=\mathcal{U}^{3}=\mathcal{H}$ and $0\in\mathrm{ran}\mathcal{Q}$, and thus, $0\in\mathrm{sri}(\mathrm{ran}\mathcal{Q}-\mathrm{ran}\mathcal{A})$, which fulfills Assumption 4-(ii). Then it follows that the scheme (3) admits any $(u,w,z)\in\mathcal{U}^{3}$ in any iterative step $k$.
A standard result in the seminal work [35] claims $z^{k}\rightharpoonup z^{\star}$ as $k\rightarrow\infty$, if $\tau>0$. This can also easily be obtained by our Theorem 5.3. The weak convergence of the solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ had been a longstanding problem until the advent of [47]. To obtain this result, [47] used some sophisticated tools, e.g., quasi-Fejér convergence, strong/weak topologies. In [51, Theorem 4.2], the author used many variants of PPA forms of DRS (especially the complicated prediction+correction framework initialized by [31, 30, 29]). Now, the next Proposition 6.1 shows that the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ can be directly obtained from (4) alone and via our Theorem 5.6. This is an improvement, compared to [51, Remark 3.1-(iii)], which once claimed that the convergences of $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ remain unclear from the PPA form of (4) only. This demonstrates the great potential of using the PPA interpretation. In addition, similar to [47], Proposition 6.1 also removes the assumption of maximal monotonicity of $\partial f+\partial g$, but without using the instrumental result of [3, Corollary 3] as [47] did.
@@
Let $\{u^{k},w^{k},z^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (3). If $\tau>0$, then $(u^{k},w^{k},z^{k})\rightharpoonup(u^{\star},w^{\star},z^{\star})$ as $k\rightarrow\infty$, where $u^{\star}$ is an optimal solution to $\min_{u\in\mathcal{U}}f(u)+g(u)$.
@@
Putting $b_{2}=b_{1}-a+c$ (indicated by the last line) into the first line, we obtain $\tau\partial g(b_{1})+b_{1}-a+c-c=\tau\partial g(b_{1})+b_{1}-a\owns 0$, leading to $b_{1}=\operatorname{prox}_{\tau g}(a)$, and thus, $b_{2}=\operatorname{prox}_{\tau g}(a)-a+c=c-\operatorname{prox}_{\tau g^{*}(\cdot/\tau)}(a)$ [22, Eq.(2.27)]. $b$ is obviously single-valued, which meets Theorem 4.2. Then, from the well-known nonexpansiveness of proximity operator (see [22, Lemma 2.4] for example), we have
@@
$=$ $\big{\|}\operatorname{prox}_{\tau g}(a)-\operatorname{prox}_{\tau g}(a^{\prime})\big{\|}+\big{\|}c-\operatorname{prox}_{\tau g^{*}(\cdot/\tau)}(a)-c^{\prime}+\operatorname{prox}_{\tau g^{*}(\cdot/\tau)}(a^{\prime})\big{\|}$